4. Portfolio risk and return

Elle holds a $10,000 portfolio that consists of four stocks. Her investment in each stock, as well as each stock’s beta, is listed in the following table:

Stock

Investment

Beta

Standard Deviation

Omni Consumer Products Co. (OCP) $3,500 0.80 15.00%
Tobotics Inc. (TI) $2,000 1.90 12.00%
Three Waters Co. (TWC) $1,500 1.20 16.00%
Makissi Corp. (MC) $3,000 0.50 28.50%

If the risk-free rate is 7% and the market risk premium is 9%, what is Elle’s portfolio’s beta and required return? Fill in the following table:

Beta

Required Return

Elle’s portfolio      
Answer #1

Total Investment = 3,500 + 2,000 + 1,500 + 3,000 = 10,000

Beta of Portfolio = Beta of OCP * Weight of OCP + Beta of TI * Weight of TI + Beta of TWC* Weight of TWC + Beta of MC* Weight of MC

= 0.80 * 3500/10000 + 1.90 * 2000/10000 + 1.2 * 1500/10000 + 0.50 * 3000/10000

= 0.99

Return of Portfolio = R: + beta(Rm - R)

= 7% + 0.99(9%)

= 15.91%

NOTE: The answer to your question has been given below/above. If there is any query regarding the answer, please ask in the comment section. If you find the answer helpful, do upvote. Help us help you.